What is the slope-intercept form of x-3y=4
Found 2 solutions by solver91311, Maths68:
Answer by solver91311(16868)



Add -x to both sides. Multiply both sides by -1/3

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by Maths68(1474)

x-3y=4
Equation of the line slope-intercept form
y=mx+b
m=slope
b=y-intercept
Rearrange the given equation accroding to the slope-intercept form
x-3y=4
-3y=-x-4
-3y=-(x+4)
Divide by -3 both sides of the equation
-3y/-3=-(x+4)/-3
y=(x+4)/3
y=x/3+4/3
y=(1/3)x+4/3
Compare the above equation with Equation of the line slope-intercept form
Slope=m=1/3
y-intercept = b = 4/3